# VGGT-Long: Chunk it, Loop it, Align it – Pushing VGGT’s Limits on Kilometer-scale Long RGB Sequences

### v1

@@ 1 Introduction
The computational and memory demands of Transformer based foundation models severely limit their scalability. Standard self-attention [39] scales quadratically with input size, and while techniques like Flash-Attention [7, 6] reduce compute complexity to linear. However, GPU memory requirement still remains prohibitive. For example, VGGT can just process 60 to 80 images on a 24 GiB RTX 4090 GPU; scaling to a KITTI Seq 00 trajectory (about 4,600 frames) would require 1,380 to 1,840 GiB, far exceeding current hardware. This bottleneck confines such models to small-scale scenes, as both memory and drift accumulation become intractable over long sequences.
Our work is inspired by recent efforts to integrate these foundation models into large-scale systems. A notable example is MASt3R-SLAM [24], which builds a sophisticatedly designed SLAM system on top of the MASt3R model. To achieve global consistency, it employs a complex backend with pose graph optimization, bundle adjustment and other heavy machinery. While powerful, such systems often entail significant engineering complexity, making them difficult to adapt to down-streaming tasks.
This raises a fundamental question: must large-scale reconstruction always equate to system-level complexity? Our philosophy diverges significantly from this trend. We advocate for a minimalist approach that unlocks the inherent potential of the foundational model itself. We posit that VGGT is already a remarkably powerful engine for large-scale 3D perception, and the primary challenge is not a lack of capability, but a lack of scalability. Instead of building another intricate system around it, we ask: can we solve the problem with the minimal overhead?
To this end, we propose VGGT-Long, a framework that extends VGGT to long sequences through a simple yet effective framework: processing the sequence in overlapping chunks, robustly aligning adjacent chunks, and correcting for drift using a high-quality loop closure module. This “chunk-and-align” paradigm avoids the need for a complex graph-based optimization backend (such as bundle adjustment [36]). It is a testament to the power of the underlying VGGT model, demonstrating that with the right strategy, its exceptional local reconstruction capabilities can be seamlessly stitched together to form a globally consistent, kilometer-scale map. Our work champions the idea that, a sufficiently powerful base model may not necessarily require a complex backend system to assist.
@@ 2 Related Work
SLAM and Visual Odometry. To overcome the scalability issue of foundation models, researchers have explored to integrate them into larger SLAM systems. Traditional SLAM systems [23, 22, 3] rely on handcrafted features and optimization, while learning-based approaches [16, 2, 34] integrate differentiable components into deep networks. However, these methods either scale poorly to long sequences or require pre-calibrated camera. MASt3R-SLAM [24] builds a sophisticated real-time SLAM framework around the MASt3R [17] without the calibration, employing complex backend machinery such as pose graph optimization and bundle adjustment to ensure global consistency.
@@ 3.1 Sequence Chunking and Local Aligning with Confidence
Given a long image sequence =$\mathcal{I}=\{\mathbf{I}_{1},\dots,\mathbf{I}_{N}\}$, we partition it into $K$ overlapping chunks. Let $L$ be the chunk size and $O$ be the overlap size. The $k$-th chunk, for $k=1,\dots,K$, is defined as the subsequence of frames indexed from $(k-1)(L-O)$ to $(k-1)(L-O)+L$. Each chunk $\mathcal{C}_{k}$ is processed independently by the VGGT model [41], which outputs a set of locally consistent camera poses and a 3D point map $\mathbf{P}_{k}\in\mathbb{R}^{H\times W\times 3}$ with the corresponding confidence values $\mathbf{c}_{k}\in\mathbb{R}^{H\times W}$.
@@
where $r_{i}^{(t)}=\|\mathbf{p}_{k}^{i}-\mathbf{S}^{(t)}\mathbf{p}_{k+1}^{i}\|_{2}$ is the residual from the previous iteration. Each weighted least-squares problem is solved efficiently in closed form using a weighted version of the Umeyama algorithm. Through this approach, we exclude low-confidence points (e.g., rapidly moving objects and sky regions) as outliers (see Fig. 4). For medium-confidence objects (e.g., slowly moving vehicles), we assign reduced alignment weights, while concentrating higher weights on high-confidence structures (e.g., buildings). In our implementation, we directly discard points with confidence values below 0.1×0.1$\times$0.1 × the median confidence of the entire chunk. The remaining low-confidence points that survive this filtering contribute minimally to the alignment process due to their attenuated weights.
@@ 3.2 Loop Detection and Loop-wise SIM(3) Aligning
Figure 4: Confidence-aware alignment suppress the influence of high-speed dynamic objects (such as vehicles) on alignment and reconstruction. It could be observed that higher-density vehicles cannot be effectively filtered out by the LiDAR, but VGGT-Long has the ability to handle this situation.
@@
The resulting 3D point map, which we can call it the “loop-centric” chunk, is then aligned with the original point maps of the corresponding chunks $\mathcal{C}_{i}$ and $\mathcal{C}_{j}$ (which were generated from temporally continuous frames). To compute the final loop-closing transformation $\mathbf{S}_{ij}$, we chain the alignments through this new chunk. Specifically, we compute the transformations from the original chunks to the loop-centric chunk and then compose them
$\mathbf{S}_{ji}=\mathbf{S}_{j,\text{loop}}\circ\mathbf{S}_{i,\text{loop}}^{-1},$ (4)
where $\mathbf{S}_{i,\text{loop}}$ and $\mathbf{S}_{j,\text{loop}}$ are the transformations that align the loop-centric chunk to the coordinate frames of chunks $\mathcal{C}_{i}$ and $\mathcal{C}_{j}$ respectively (described in Sec. 3.1). The composition $\mathbf{S}_{j,\text{loop}}\circ\mathbf{S}_{i,\text{loop}}^{-1}$ effectively chains the transformations to compute the direct alignment from chunk $\mathcal{C}_{i}$ to $\mathcal{C}_{j}$ (i.e., $i\rightarrow\text{loop}\rightarrow j$). This provides a robust geometric constraint for the global optimization by bridging the two distant chunks through a shared, high-quality local reconstruction.
@@ 4 Experiments
Classic ORB-SLAM2 (w/o LC) [22] Required Sparse 69.727 26.480 40.65 502.20 47.82 [0.94] 1.30 29.95 40.82 16.04 (43.09) (38.77) [5.42]
ORB-SLAM2 (w/ LC) [22] Required Sparse 54.816 [9.464] [6.03] 508.34 [14.76] (1.02) 1.57 [4.04] (11.16) (2.19) [38.85] [8.39] (6.63)
@@
DPV-SLAM++ [20] Required Sparse (25.749) 27.138 (8.30) 11.86 39.64 2.50 (0.78) 5.74 11.60 [1.52] 110.90 76.70 13.70
@@
VGGT-Long (Ours) No Need Dense 27.637 (18.284) 8.67 121.17 32.08 6.12 4.23 8.31 [5.34] 4.63 53.10 41.99 18.37
Table 1: Camera Tracking Results (ATE RMSE [m] $\downarrow$) on the KITTI Dataset. [LC] denotes loop closure. Since the Seq. 01 is a high-speed sequence, its movement pattern is significantly different from those of other sequences. [Avg.] shows the mean ATE excluding Seq 01. VGGT-Long achieves relatively good tracking accuracy without the input of camera intrinsic matrix. [OOM] is short for CUDA Out-Of-Memory on a single RTX 4090. [TL] is short for Tracking Lost. See Fig. 7 for trajectory visual results. Color defination: [first], (second), third.
@@
Table 2: Camera Tracking Results (ATE RMSE [m] $\downarrow$) on the Waymo Open Dataset. Color defination: [first], (second).
@@
Table 3: Point Map Estimation Results on the Waymo Open Dataset. The GT point cloud is collected by LiDAR, which on vehicles is typically scanning at a height below human eye level, the GT point cloud’s coverage is actually narrower than the observation range of RGB cameras. As can be seen in Segment 371159869, the LiDAR-captured GT point cloud fails to perceive the 3D structure of the overpass, whereas almost all visual-based methods can, to varying degrees, detect and reconstruct this structure. Therefore, the numerical values in this table should be taken as reference only, and readers are encouraged to refer to Fig. 9 to Fig. 12 for a more intuitive visual comparison between methods. However, whether by visual comparison or the reference numerical values, it is evident that VGGT-Long achieves high reconstruction performance than other methods. Color defination: [first].
@@
Table 4: Camera Tracking Results (ATE RMSE [m] $\downarrow$) on the Virtual KITTI Dataset. Color defination: [first], (second).
We evaluate VGGT-Long on three datasets — KITTI Dataset Odometry Track [13], Waymo Open Dataset (v1.4.1) [31], and Virtual KITTI Dataset (v1.3.1) [11] to assess its performance in large-scale monocular 3D reconstruction. Our experiments are conducted on the computer with Ubuntu 22.04, and equipped with 12 Intel Xeon Gold 6128 3.40 GHz CPUs, 67GiB of RAM, and an NVIDIA RTX 4090 GPU with 24GiB of VRAM.
@@ 4.2 Experiments Settings
For the calculation of reconstruction performance metrics, to avoid interference from outlier point clouds, we uniformly retain points with confidence values greater than 0.75×0.75$\times$0.75 × the average confidence for Transformer-based 3D methods. As for models like DROID-SLAM, which do not output confidence values, we retain points with depth values larger than 0.75×0.75$\times$0.75 × the average inverse depth as the filtered point cloud.
Following the VGGT’s configuration, we downsample input images to 518-pixel width while the hight maintaining aspect ratio during inference. Due to the more “elongated” image aspect ratios in KITTI and Virtual KITTI datasets, we set the chunk size to 75 frames, while adopting 60 frames for Waymo (nearly saturating the whole capacity of RTX 4090 24GiB VRAM). Our inference pipeline first processes all images through VPR, then releases the VPR model’s memory occupancy to allocate sufficient GPU memory resources for VGGT.
@@ 4.4 KITTI, Waymo & Virtual KITTI
VGGT-Long outperforms the learning-based methods such as DROID-SLAM and DPVO. Unlike ORB-SLAM2 or LDSO, our method does not rely on calibrated camera intrinsics and still maintains competitive accuracy. Notably, VGGT-Long runs successfully on all sequences, while foundation models like Fast3R, CUT3R, and VGGT fail due to memory overflow. This verifies that our chunk-and-align framework effectively extends VGGT to long sequences without sacrificing scalability.

### v2

# VGGT-Long: Chunk it, Loop it, Align it, Pushing VGGT’s Limits on Kilometer-scale Long RGB Sequences

@@ 1 Introduction
The computational and memory demands of Transformer based foundation models severely limit their scalability. Standard self-attention [38] scales quadratically with input size, and while techniques like Flash-Attention [4, 5] reduce compute complexity to linear. However, GPU memory requirement still remains prohibitive. For example, VGGT can just process 60 to 80 images on a 24 GiB RTX 4090 GPU scaling to a KITTI Seq 00 trajectory (about 4,600 frames) would require unreachable GPU memory requirement, far exceeding current hardware. This bottleneck confines such models to small-scale scenes, as both memory and drift accumulation become intractable over long sequences.
Our work is inspired by recent efforts to integrate foundation models into large-scale systems. A notable example is MASt3R-SLAM [23], which builds SLAM system on top of the MASt3R [15] model. To achieve global consistency, it employs pose graph optimization and bundle adjustment within its backend, which are standard components in modern SLAM systems.
This raises a fundamental question: must large-scale reconstruction always equate to system-level complexity? Our philosophy diverges significantly from this trend. We advocate for a minimalist approach that unlocks the inherent potential of the foundational model itself. We posit that VGGT is already a remarkably powerful engine for large-scale 3D perception, and the primary challenge is not a lack of capability, but a lack of scalability. Instead of building another full system around it, we ask: can we solve the problem with the minimal overhead?
To this end, we propose VGGT-Long, a framework that extends VGGT to long sequences through a simple yet effective framework that is processing the sequence in overlapping chunks, robustly aligning adjacent chunks, and correcting for drift using a high-quality loop closure module. This “chunk-and-align” paradigm avoids the need for a graph-based optimization backend (such as bundle adjustment [35]). It is a testament to the power of the underlying VGGT model, demonstrating that with the right strategy, its exceptional local reconstruction capabilities can be seamlessly stitched together to form a globally consistent, kilometer-scale map. Our work champions the idea that, a sufficiently powerful base model may not necessarily require a system-level backend to assist.
@@ 2 Related Work
SLAM and Visual Odometry. To overcome the scalability issue of foundation models, researchers have explored to integrate them into larger SLAM systems. Traditional SLAM systems [21, 22] rely on handcrafted features and optimization, while learning-based approaches [14, 33] integrate differentiable components into deep networks. However, these methods either scale poorly to long sequences or require pre-calibrated camera. MASt3R-SLAM [23] builds a sophisticated real-time SLAM framework around the MASt3R [15] without the calibration, employing complex backend machinery such as pose graph optimization and bundle adjustment to ensure global consistency. Concurrent work VGGT-SLAM [19] introduces a complete SLAM system that aligns submaps via SL(4)-based factor graph optimization for accurate indoor reconstruction. In contrast, our VGGT-Long targets large-scale outdoor scenes using a lightweight pipeline with Sim(3) transformations, prioritizing simplicity and scalability over a full SLAM framework.
@@ 3.1 Sequence Chunking and Local Aligning with Confidence
Given a long image sequence $\mathcal{I}=\{\mathbf{I}_{1},\dots,\mathbf{I}_{N}\}$, we partition it into $K$ overlapping chunks. Let $L$ be the chunk size and $O$ be the overlap size. The $k$-th chunk, for $k=1,\dots,K$, is defined as the subsequence of frames indexed from (k1)(LO)(k-1)(L-O) to (k1)(LO)+L(k-1)(L-O)+L. Each chunk $\mathcal{C}_{k}$ is processed independently by the VGGT model [39], which outputs a set of locally consistent camera poses and a 3D point map $\mathbf{P}_{k}\in\mathbb{R}^{H\times W\times 3}$ with the corresponding confidence values $\mathbf{c}_{k}\in\mathbb{R}^{H\times W}$.
@@
where $r_{i}^{(t)}=\|\mathbf{p}_{k}^{i}-\mathbf{S}^{(t)}\mathbf{p}_{k+1}^{i}\|_{2}$ is the residual from the previous iteration. Each weighted least-squares problem is solved efficiently in closed form using a weighted version of the Umeyama algorithm. Through this approach, we exclude low-confidence points (e.g., rapidly moving objects and sky regions) as outliers (see Fig. 4). For medium-confidence objects (e.g., slowly moving vehicles), we assign reduced alignment weights, while concentrating higher weights on high-confidence structures (e.g., buildings). In our implementation, we directly discard points with confidence values below $0.1\times$ the median confidence of the entire chunk. The remaining low-confidence points that survive this filtering contribute minimally to the alignment process due to their attenuated weights.
@@ 3.2 Loop Detection and Loop-wise SIM(3) Aligning
Figure 4: Confidence-aware alignment suppresses the influence of high-speed dynamic objects (such as vehicles) on alignment and reconstruction. It could be observed that higher-density vehicles cannot be effectively filtered out by the LiDAR, but VGGT-Long has the ability to handle this situation.
@@
Figure 5: Loop-wise Sim(3) Alignment
The resulting 3D point map, which we can call it the “loop-centric” chunk (see Figure 5), is then aligned with the original point maps of the corresponding chunks $\mathcal{C}_{i}$ and $\mathcal{C}_{j}$ (which were generated from temporally continuous frames). To compute the final loop-closing transformation $\mathbf{S}_{ij}$, we chain the alignments through this new chunk. Specifically, we compute the transformations from the original chunks to the loop-centric chunk and then compose them
$\mathbf{S}_{ij}=\mathbf{S}_{i,\text{loop}}\circ\mathbf{S}_{j,\text{loop}}^{-1},$ (4)
where $\mathbf{S}_{i,\text{loop}}$ and $\mathbf{S}_{j,\text{loop}}$ are the transformations that align the loop-centric chunk to the coordinate frames of chunks $\mathcal{C}_{i}$ and $\mathcal{C}_{j}$ respectively (described in Sec. 3.1). The composition $\mathbf{S}_{i,\text{loop}}\circ\mathbf{S}_{j,\text{loop}}^{-1}$ effectively chains the transformations to compute the direct alignment from chunk $\mathcal{C}_{i}$ to $\mathcal{C}_{j}$ (i.e., $i\rightarrow\text{loop}\rightarrow j$). This provides a robust geometric constraint for the global optimization by bridging the two distant chunks through a shared, high-quality local reconstruction.
@@ 4 Experiments
Classic ORB-SLAM2 (w/o LC) [22] Required Sparse 69.727 26.480 40.65 502.20 47.82 [0.94] 1.30 29.95 40.82 16.04 (43.09) 38.77 [5.42]
ORB-SLAM2 (w/ LC) [22] Required Sparse 54.816 [9.464] [6.03] 508.34 [14.76] (1.02) 1.57 [4.04] 11.16 (2.19) [38.85] [8.39] (6.63)
@@
DPV-SLAM++ [18] Required Sparse 25.749 27.138 8.30 11.86 39.64 2.50 (0.78) 5.74 11.60 [1.52] 110.90 76.70 13.70
@@
VGGT-Long (Chunk Size=30) No Need Dense 44.713 39.564 9.06 96.20 99.95 19.76 11.36 10.20 10.04 4.00 139.79 50.53 40.94
VGGT-Long (Chunk Size=60) No Need Dense 26.358 (19.298) (8.06) 96.96 34.16 6.83 4.16 9.15 [4.68] (2.68) 63.15 32.24 27.87
VGGT-Long (Chunk Size=90) No Need Dense (22.718) 20.938 11.97 40.51 49.85 5.41 2.86 9.88 (6.07) 3.47 66.27 32.27 21.33
VGGT-Long (Chunk Size=120) No Need Dense 25.597 22.814 16.13 53.43 51.98 4.37 2.15 12.69 11.33 3.603 70.29 34.55 21.05
Table 1: Camera Tracking Results (ATE RMSE [m] $\downarrow$) on the KITTI Dataset Odometry Track. The overlap was set to half of the chunk size. [LC] denotes loop closure. Since the Seq. 01 is a high-speed sequence, its movement pattern is significantly different from those of other sequences. [Avg.] shows the mean ATE excluding Seq 01. VGGT-Long achieves relatively good tracking accuracy without the input of camera intrinsic matrix. [OOM] is short for CUDA Out-Of-Memory on a single RTX 4090. [TL] is short for Tracking Lost. See Fig. 8 for trajectory visual results. Color definition: [first], (second), third.
@@
Table 2: Camera Tracking Results (ATE RMSE [m] $\downarrow$) on the Waymo Open Dataset. Color definition: [first], (second).
@@
Table 3: Point Map Estimation Results on the Waymo Open Dataset. The GT point cloud is collected by LiDAR, which on vehicles is typically scanning at a height below human eye level, the GT point cloud’s coverage is actually narrower than the observation range of RGB cameras. As can be seen in Segment 371159869, the LiDAR-captured GT point cloud fails to perceive the 3D structure of the overpass, whereas almost all visual-based methods can, to varying degrees, detect and reconstruct this structure. Therefore, the numerical values in this table should be taken as reference only, and readers are encouraged to refer to Fig. 10 to Fig. 13 for a more intuitive visual comparison between methods. However, whether by visual comparison or the reference numerical values, it is evident that VGGT-Long achieves high reconstruction performance than other methods. Color definition: [first].
@@
Table 4: Camera Tracking Results (ATE RMSE [m] $\downarrow$) on the Virtual KITTI Dataset. Color definition: [first], (second).
We evaluate VGGT-Long on three datasets, KITTI Dataset Odometry Track [11], Waymo Open Dataset (v1.4.1) [30], and Virtual KITTI Dataset (v1.3.1) [9] to assess its performance in large-scale monocular 3D reconstruction. Our experiments are conducted on the computer with Ubuntu 22.04, and equipped with 12 Intel Xeon Gold 6128 3.40 GHz CPUs, 67GiB of RAM. In this paper, most experiments were conducted using an NVIDIA RTX 4090 GPU with 24 GiB of VRAM, while for experiments with a chunk size of 90 or larger, we used an NVIDIA L20 GPU with 48 GiB of VRAM.
@@ 4.2 Experiments Settings
For the calculation of reconstruction performance metrics, to avoid interference from outlier point clouds, we uniformly retain points with confidence values greater than $0.75\times$ the average confidence for Transformer-based 3D methods. As for models like DROID-SLAM, which do not output confidence values, we retain points with depth values larger than $0.75\times$ the average inverse depth as the filtered point cloud.
Following the VGGT’s configuration, we downsample input images to 518-pixel width while the hight maintaining aspect ratio during inference. Our inference pipeline first processes all images through VPR, then releases the VPR model’s memory occupancy to allocate sufficient GPU memory resources for VGGT.
@@ 4.4 KITTI, Waymo & Virtual KITTI
VGGT-Long outperforms the learning-based methods such as DROID-SLAM and DPVO. Unlike ORB-SLAM2 or LDSO, our method does not rely on calibrated camera intrinsics and still maintains competitive accuracy. Notably, VGGT-Long runs successfully on all sequences, while foundation models like Fast3R, CUT3R, and VGGT fail due to memory overflow. This verifies that our chunk-and-align framework effectively extends VGGT to long sequences without sacrificing scalability. When running MASt3R-SLAM on KITTI Odometry, we observe that tracking stalled after about 100 frames. New frames cannot be selected as keyframes, causing the mapping part to stop updating even though the process continues running. This phenomenon indicates that MASt3R-SLAM undergoes tracking lost. Tracking lost is defined when the system fails to register new frames for a prolonged period, leading to no further updates in the map and trajectory. Although Figure 6 shows a better reconstruction correction and Table 1 demonstrates better tracking performance, we can still observe that there are still some parts of the central intersection that are not aligned. This is because there is no explicit loop closure constraint for the input image in this area.